# On Edge Coloring of Multigraphs

### v4

@@ 3 Extented Tashkinov Tree (ETT)
Let $n^{\prime}$ be the smallest index such that $v_{1}\in T_{n^{\prime}}$ and $T_{n^{\prime}}$ is RE finished, and let $h^{\prime}$ be the initial index of $T_{n^{\prime}}$. So $n^{\prime}\leq h$, and $v_{1}\notin T_{h^{\prime}}$. Let $\alpha^{\prime}\in\overline{\varphi}(v_{1})=\overline{\sigma}(v_{1})$ and $\tau\in\overline{\varphi}(x_{e})$ with $\tau\neq\alpha^{\prime}$ and $\tau\neq\gamma_{h^{\prime}}$; we have such choices because $T_{n+1}$ is elementary under $\varphi$, $v_{1}$ is not the first vertex in $T_{n+1}$ along $\prec$, and $|\overline{\varphi}(x_{e})|\geq 2$. So $\alpha^{\prime},\tau\notin\{\delta_{h},\eta\}$ and $\alpha^{\prime}\neq\gamma$ by the choice of $\alpha^{\prime},\tau$, (2), $\delta_{h}\notin\overline{\varphi}(T_{n+1})$, and the elementariness of $T_{n+1}$ under $\varphi$. Moreover, under $\sigma$ and $\varphi$, $T_{h}$ is closed for both $\alpha^{\prime}$ and $\gamma_{h}$, and $T_{n+1}$ and $T_{h}$ are closed for both $\gamma$ and $\tau$ ($\gamma=\tau$ may happen). Let $\mu=\sigma/(G-T_{h},\gamma,\tau)$. Let $\mu_{1}=\mu/(G-T_{h},\alpha^{\prime},\gamma_{h})$ if $\tau\neq\gamma_{h}$, and $\mu_{1}=\mu/(G-T_{h},\alpha^{\prime},\gamma)$ otherwise. Furthermore, because $P$ is disjoint from $T_{n}$, following (8), we have the following under $\mu_{1}$:
@@
(9) $\mu_{1}$ is $(T_{n^{\prime}},D_{n^{\prime}},\varphi)$-stable, $T_{n+1}$ is elementary under $\mu_{1}$, $P_{ex}$ is a $(\alpha^{\prime},\delta_{h})$-exit path of both $T_{n^{\prime}}$ and $T_{h}$ with $\alpha^{\prime}\in\overline{\mu}_{2}(v_{1})$, $D$ is an $(\alpha^{\prime},\delta_{h})$-ear of $T_{h}$, $\mu_{1}(P)=\{\tau,\eta,\delta_{h}\}$, all the $\eta$-colored edges of $P$ are not in $G[V(T_{n+1})]$, all the $\tau$ colored edges of $P$ are in $G[V(T_{n+1})]$, $\delta_{h}\in\overline{\mu}_{1}(b_{j})$, two ends of $P$ miss either $\delta_{h}$ or $\eta$, exactly one boundary edge of $T_{n+1}$ is colored by $\eta$ which is $f_{j}$ under $\mu_{1}$, and $T_{n+1}$ is closed for $\tau$ under $\mu_{1}$.
@@
Because $T_{n+1}$ is elementary under $\mu_{1}$, exactly one boundary edge of $T_{n+1}$ is colored by $\eta$ which is $f_{j}$ under $\mu_{1}$, and $T_{n+1}$ is closed for $\tau$ under $\mu_{1}$, we see that there is exactly one $(\tau,\eta)$-path $P_{x_{e}}(\eta,\tau,\mu_{1})$ intersecting $T_{n+1}$ under $\mu_{1}$ with $(\tau,\eta)$-exit $b_{j}$. Moreover, the subpath of $P_{x_{e}}(\eta,\tau,\mu_{1})$ from $x_{e}$ to $b_{j}$ is contained in $G[V(T_{n+1})]$, and all the other $(\tau,\eta)$-chains intersecting $T_{n+1}$ are cycles also contained in $G[V(T_{n+1})]$. So following (9), all the $\eta$-colored edges of $P$ are contained in $(\tau,\eta)$-chains not intersecting $T_{n+1}$. Consequently, following (9), because all the $\tau$ colored of $P$ are in $G[V(T_{n+1})]$ and no $\eta$-colored edges of $P$ are in $G[V(T_{n+1})]$, the coloring $\mu_{2}$ obtained from $\mu_{1}$ by switching all the $\eta$-colored edges of $P$ not in $P_{x_{e}}(\eta,\tau,\mu_{1})$ to $\tau$ by Kempe changes, satisfies the following:
@@
    $\mu_{2}$ is $(T_{n^{\prime}},D_{n^{\prime}},\varphi)$-stable, $P_{ex}$ is a $(\alpha^{\prime},\delta_{h})$-exit path of both $T_{n^{\prime}}$ and $T_{h}$ with $\alpha^{\prime}\in\overline{\mu}_{2}(v_{1})$, and $D$ is an $(\alpha^{\prime},\delta_{h})$-ear of $T_{h}$ under $\mu_{2}$.
@@
    All the $\eta$-colored edges of $P$ belong to $P_{x_{e}}(\eta,\tau,\mu_{2})=P_{x_{e}}(\eta,\tau,\mu_{1})$ containing the vertex $b_{j}\in P$ with one end $x_{e}\in T_{n^{\prime}}(v_{1})$ missing $\tau$ and the other end $r^{\prime}$ is outside of $T_{n^{\prime}}$, and all of these $\eta$-colored edges of $P$ are in the subpath from $b_{j}$ to $r^{\prime}$ of $P_{x_{e}}(\eta,\tau,\mu_{2})$, and this subpath does not intersect $T_{n^{\prime}}$.
Because $T_{n^{\prime}}$ is $RE$ finished, following (A2) on $T_{n^{\prime}}$ and (o), we see that $T_{n^{\prime}}$ is an ETT with $(n^{\prime}-1)$-rungs under $\mu_{2}$, and $v_{2}$ is a root of an $(\alpha^{\prime},\delta_{h})$-ear $D^{\prime}$ of $T_{n^{\prime}}$. Since $v_{2}$ is a root of the $(\alpha^{\prime},\delta_{h})$-ear $D$ of $T_{h}$ with $h\geq n^{\prime}$, $D$ is a subpath of $D^{\prime}$, and $P$ shares edges with both $P_{ex}$ and $D^{\prime}$ by (i). Moreover, since $\tau\in\overline{\varphi}(x_{e})$, we have $\tau\neq\delta_{h^{\prime}}$. So $\tau\notin S_{h^{\prime}}$. Thus with (o), (i), (ii), (iii) and (iv), we have a contradiction to the later introduced technical Lemma 4.6 under $\mu_{2}$ with $T_{n^{\prime}}$ ($n^{\prime}\leq n$) in place of $T_{n+1}$, $D^{\prime}$ in place of $D$, $\alpha^{\prime}$ in place of $\alpha$, $\delta_{h}$ in place of $\delta$, $x_{e}$ in place of $r$, $v_{1}$ in place of $x$, $v_{2}$ in place of $y$, and $b_{j}$ in place of $w$. ∎
@@
(A2.5) Let $n$ be a nonnegative integer. Suppose (A1) and (A2) hold for all ETTs with at most $n$ rungs, and $(A2.5)$ holds for all ETTs with at most $(n-1)$ rungs. Let $T_{n+1}$ be a closed ETT with ladder $T_{0}\subset\dots\subset T_{n}\subset T_{n+1}$. Suppose $T_{n+1}$ could be extended further to $T_{n+2}$ with $F_{n+1}=\{f_{n+1}\}$, $S_{n+1}=\{\gamma_{n+1},\delta_{n+1}\}$, and initial index $h$. If $\varphi^{\prime}$ is a $(T_{n+1},D_{n+1},\varphi)$-stable coloring such that $T_{h}$ has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\varphi^{\prime}$ as under $\varphi$, then $\varphi^{\prime}$ is a $T_{n+1}+f_{n+1}$ mod $\varphi$ coloring.
@@
Let $T_{n+1}$, $F_{n+1}$, $S_{n+1}$, $h$, and $\varphi^{\prime}$ be as described above. We first consider the case when extension type $\Theta_{n+1}=IE$. Then $T_{n+1}$ is $RE$ finished, $h=n+1$, $\delta_{n+1}$ is a defective color of $T_{n+1}$, and $f_{n}$ belongs to a $(\gamma_{n+1},\delta_{n+1})$-ear of $T_{n+1}$ under $\varphi$. Moreover, since $T_{n+1}$ is elementary by (A1), $T_{n+1}$ has exactly one $(\gamma_{n+1},\delta_{n+1})$-exit vertex under $\varphi$ by (A2). Since $\varphi^{\prime}$ is $(T_{n+1},D_{n+1},\varphi)$-stable, $T_{n+1}$ is closed under $\varphi^{\prime}$ and $\varphi^{\prime}$ is $(T_{n+1},D_{n},\varphi)$-stable. So $\varphi^{\prime}$ is a $T_{n+1}$ mod $\varphi$ coloring by (A2), and therefore (1) $T_{n+1}$ is a closed ETT under $\varphi^{\prime}$ with the same ladder, $F_{i}$, $S_{i}$, and $\Theta_{i}$ for $0\leq i\leq n$, and $T_{n+1}$ is still $RE$ finished under $\varphi^{\prime}$. Recall that $f_{n+1}$ belongs to a $(\gamma_{n+1},\delta_{n+1})$-ear of $T_{n+1}$ under $\varphi$. Since $\varphi^{\prime}$ is a $(T_{n+1},D_{n+1},\varphi)$-stable coloring such that $T_{n+1}$ has the same $(\gamma_{n+1},\delta_{n+1})$-exit vertex under $\varphi^{\prime}$ as $\varphi$, the connecting edge $f_{n+1}$ still belongs to a $(\gamma_{n+1},\delta_{n+1})$-ear of $T_{n+1}$ under $\varphi^{\prime}$. Since $T_{n+1}$ is still $RE$ finished under $\varphi^{\prime}$ by (1), we may still choose $F_{n+1}=\{f_{n+1}\}$ and $S_{n+1}=\{\gamma_{n+1},\delta_{n+1}\}$ to extend $T_{n+1}$ by an $IE$ extension under $\varphi^{\prime}$, so the coloring $\varphi^{\prime}$ is a $T_{n+1}+f_{n+1}$ mod $\varphi$ coloring following (1). We then consider the case when extension type $\Theta_{n+1}=RE$ with initial index $h\leq n$. In this case, $n\geq 1$ and $h\geq 1$. Since $T_{h}$ is elementary by (A1), $T_{h}$ has exactly one $(\gamma_{h},\delta_{h})$-exit vertex under $\varphi$ by (A2). By the assumption on $\varphi^{\prime}$, the ETT $T_{h}$ has the same unique $(\gamma_{h},\delta_{h})$-exit vertex under $\varphi^{\prime}$ as under $\varphi$. Note that in this case $\varphi^{\prime}$ is also a $(T_{n},D_{n},\varphi)$-stable coloring, we see that $\varphi^{\prime}$ is a $T_{n}+f_{n}$ mod $\varphi$ coloring by (A2.5) with $T_{n}$ that has $(n-1)$-rungs in place of $T_{n+1}$. As $\varphi^{\prime}$ is also $(T_{n+1},D_{n+1},\varphi)$-stable, (2) $T_{n+1}$ is a closed ETT under $\varphi^{\prime}$ with the same ladder, $F_{i}$, $S_{i}$, and $\Theta_{i}$ for $0\leq i\leq n$. Furthermore, $f_{n+1}$ is still colored by $\gamma_{h}=\gamma_{n+1}$. Because $\Theta_{n+1}=RE$, the connecting edge $f_{n+1}$ belongs to a $(\gamma_{h},\delta_{h})$-ear $O$ of $T_{h}$ such that $O$ contains a path $L$ connecting $f_{n+1}$ and $V(T_{h})$ with $V(L)\subseteq V(T_{n+1})$ under $\varphi$. Since $\varphi^{\prime}$ is $(T_{n+1},D_{n+1},\varphi)$-stable, $L$ remains a path contained entirely in either a $(\gamma_{h},\delta_{h})$-ear or a $(\gamma_{h},\delta_{h})$-exit path of $T_{h}$ under $\varphi^{\prime}$. Since $T_{h}$ has the same unique $(\gamma_{h},\delta_{h})$-exit vertex under $\varphi^{\prime}$ as under $\varphi$, the path $L$ must be contained entirely in a $(\gamma_{h},\delta_{h})$-ear $O^{*}$ of $T_{h}$ under $\varphi^{\prime}$ that connects $f_{n+1}$ to $V(T_{h})$ (here $O^{*}$ might be different from $O$). So we may still choose $F_{n+1}=\{f_{n+1}\}$ and $S_{n+1}=\{\gamma_{n+1},\delta_{n+1}\}$ to extend $T_{n+1}$ by an $RE$ extension under $\varphi^{\prime}$, and therefore the coloring $\varphi^{\prime}$ is a $T_{n+1}+f_{n+1}$ mod $\varphi$ coloring following (2). ∎
@@
(A3) Let $n$ be a nonnegative integer. Suppose (A1), (A2), and (A2.5) hold for all ETTs with at most $n$ rungs, and (A3) holds for all ETTs with at most $(n-1)$ rungs. Let $T_{n+1}$ be a closed ETT with ladder $T_{0}\subset\dots\subset T_{n}\subset T_{n+1}$. Assume $T_{n+1}$ is $RE$ finished, $\delta$ is a defective color of $T_{n+1}$ under $\varphi$, and $\gamma\in\overline{\varphi}(T_{n+1})$. Let $\alpha,\beta$ be two colors such that $\alpha\in\overline{\varphi}(T_{n+1})\setminus\{\gamma\}$ and $\beta$ is arbitrary. For any $(\alpha,\beta)$-path $P$, if $P$ does not intersect $T_{n+1}$, then $T_{n+1}$ has the same $(\gamma,\delta)$-exit vertex under $\varphi/P$ as $\varphi$.

### v5

@@ 3 Extented Tashkinov Tree (ETT)
Let $n^{\prime}$ be the smallest index such that $v_{1}\in T_{n^{\prime}}$ and $T_{n^{\prime}}$ is $\operatorname{RE}$ finished, and let $h^{\prime}$ be the initial index of $T_{n^{\prime}}$. So $n^{\prime}\leq h$, and $v_{1}\notin T_{h^{\prime}}$. Let $\alpha^{\prime}\in\overline{\varphi}(v_{1})=\overline{\sigma}(v_{1})$ and $\tau\in\overline{\varphi}(x_{e})$ with $\tau\neq\alpha^{\prime}$ and $\tau\neq\gamma_{h^{\prime}}$; we have such choices because $T_{n+1}$ is elementary under $\varphi$, $v_{1}$ is not the first vertex in $T_{n+1}$ along $\prec$, and $|\overline{\varphi}(x_{e})|\geq 2$. So $\alpha^{\prime},\tau\notin\{\delta_{h},\eta\}$ and $\alpha^{\prime}\neq\gamma$ by the choice of $\alpha^{\prime},\tau$, (2), $\delta_{h}\notin\overline{\varphi}(T_{n+1})$, and the elementariness of $T_{n+1}$ under $\varphi$. Moreover, under $\sigma$ and $\varphi$, $T_{h}$ is closed for both $\alpha^{\prime}$ and $\gamma_{h}$, and $T_{n+1}$ and $T_{h}$ are closed for both $\gamma$ and $\tau$ ($\gamma=\tau$ may happen). Let $\mu=\sigma/(G-T_{h},\gamma,\tau)$. Let $\mu_{1}=\mu/(G-T_{h},\alpha^{\prime},\gamma_{h})$ if $\tau\neq\gamma_{h}$, and $\mu_{1}=\mu/(G-T_{h},\alpha^{\prime},\gamma)$ otherwise. Furthermore, because $P$ is disjoint from $T_{n}$, following (8), we have the following under $\mu_{1}$:
@@
(9) $\mu_{1}$ is $(T_{n^{\prime}},D_{n^{\prime}},\varphi)$-stable, $T_{n+1}$ is elementary under $\mu_{1}$, $P_{ex}$ is an $(\alpha^{\prime},\delta_{h})$-exit path of both $T_{n^{\prime}}$ and $T_{h}$ with $\alpha^{\prime}\in\overline{\mu}_{2}(v_{1})$, $D$ is an $(\alpha^{\prime},\delta_{h})$-ear of $T_{h}$, $\mu_{1}(P)=\{\tau,\eta,\delta_{h}\}$, all the $\eta$ colored edges of $P$ are not in $G[V(T_{n+1})]$, all the $\tau$ colored edges of $P$ are in $G[V(T_{n+1})]$, $\delta_{h}\in\overline{\mu}_{1}(b_{j})$, two ends of $P$ miss either $\delta_{h}$ or $\eta$, exactly one boundary edge of $T_{n+1}$ is colored by $\eta$ which is $f_{j}$ under $\mu_{1}$, and $T_{n+1}$ is closed for $\tau$ under $\mu_{1}$.
@@
Because $T_{n+1}$ is elementary under $\mu_{1}$, exactly one boundary edge of $T_{n+1}$ is colored by $\eta$ which is $f_{j}$ under $\mu_{1}$, and $T_{n+1}$ is closed for $\tau$ under $\mu_{1}$, we see that there is exactly one $(\tau,\eta)$-path $P_{x_{e}}(\eta,\tau,\mu_{1})$ intersecting $T_{n+1}$ under $\mu_{1}$ with $(\tau,\eta)$-exit $b_{j}$. Moreover, the subpath of $P_{x_{e}}(\eta,\tau,\mu_{1})$ from $x_{e}$ to $b_{j}$ is contained in $G[V(T_{n+1})]$, and all the other $(\tau,\eta)$-chains intersecting $T_{n+1}$ are cycles also contained in $G[V(T_{n+1})]$. So following (9), all the $\eta$ colored edges of $P$ are contained either in $(\tau,\eta)$-chains not intersecting $T_{n+1}$ or in $P_{x_{e}}(\eta,\tau,\mu_{1})$. Consequently, following (9), because all the $\tau$ colored of $P$ are in $G[V(T_{n+1})]$ and no $\eta$ colored edges of $P$ are in $G[V(T_{n+1})]$, the coloring $\mu_{2}$ obtained from $\mu_{1}$ by switching all the $\eta$ colored edges of $P$ not in $P_{x_{e}}(\eta,\tau,\mu_{1})$ to $\tau$ by Kempe changes, satisfies the following:
@@
    $\mu_{2}$ is $(T_{n^{\prime}},D_{n^{\prime}},\varphi)$-stable, $P_{ex}$ is an $(\alpha^{\prime},\delta_{h})$-exit path of both $T_{n^{\prime}}$ and $T_{h}$ with $\alpha^{\prime}\in\overline{\mu}_{2}(v_{1})$, and $D$ is an $(\alpha^{\prime},\delta_{h})$-ear of $T_{h}$ under $\mu_{2}$.
@@
    All the $\eta$ colored edges of $P$ belong to $P_{x_{e}}(\eta,\tau,\mu_{2})=P_{x_{e}}(\eta,\tau,\mu_{1})$ containing the vertex $b_{j}\in V(P)$ with one end $x_{e}\in T_{n^{\prime}}(v_{1})$ missing $\tau$ and the other end $r^{\prime}$ is outside of $T_{n^{\prime}}$, and all of these $\eta$ colored edges of $P$ are in the subpath from $b_{j}$ to $r^{\prime}$ of $P_{x_{e}}(\eta,\tau,\mu_{2})$, and this subpath does not intersect $T_{n^{\prime}}$.
Because $T_{n^{\prime}}$ is $\operatorname{RE}$ finished, following (A2) on $T_{n^{\prime}}$ and (o), we see that $T_{n^{\prime}}$ is an ETT with $(n^{\prime}-1)$-rungs under $\mu_{2}$, and $v_{2}$ is a root of an $(\alpha^{\prime},\delta_{h})$-ear $D^{\prime}$ of $T_{n^{\prime}}$. Since $v_{2}$ is a root of the $(\alpha^{\prime},\delta_{h})$-ear $D$ of $T_{h}$ with $h\geq n^{\prime}$, $D$ is a subpath of $D^{\prime}$, and $P$ shares edges with both $P_{ex}$ and $D^{\prime}$ by (i). Moreover, since $\tau\in\overline{\varphi}(x_{e})$, we have $\tau\neq\delta_{h^{\prime}}$. So $\tau\notin S_{h^{\prime}}$. Thus with (o), (i), (ii), (iii) and (iv), we have a contradiction to the later introduced technical Lemma 4.6 under $\mu_{2}$ with $T_{n^{\prime}}$ ($n^{\prime}\leq n$) in place of $T_{n+1}$, $D^{\prime}$ in place of $D$, $\alpha^{\prime}$ in place of $\alpha$, $\delta_{h}$ in place of $\delta$, $x_{e}$ in place of $r$, $v_{1}$ in place of $x$, $v_{2}$ in place of $y$, and $b_{j}$ in place of $w$. ∎
Remark 2.
Because almost all properties remain valid when we switch $\alpha$ and $\gamma$ for edges in $G-V(T_{n+1})$, for two colors $\alpha$ and $\gamma$ missing at vertices of $T_{n+1}$, it is tempting to combine Lemmas 4.34.7 into a single lemma and apply it directly in the proof of Lemma 4.1. Indeed, Lemma 3.2 in [4] is analogous to our Lemmas 4.34.7. However, their proof is incomplete due to the absence of the key techniques developed in Lemma 4.6.
Specifically, in the second paragraph after Claim 1 of their Lemma 3.2, they define $\varphi^{*}=\varphi/(G-U,\alpha_{1},\alpha)$. They then claim that $R$, an $(\alpha_{1},\alpha_{2})$-chain under $\varphi$ which is an entire $(\alpha_{1},\alpha_{2})$-path (they define $(\alpha,\beta)$-chain as an entire $(\alpha,\beta)$-component that is a path), becomes an entire $(\alpha,\alpha_{2})$-path under $\varphi^{*}$. This claim does not hold when $R$ and $U$ share an endpoint $v$ of $R$ with $\alpha_{1}\in\overline{\varphi}(v)$ and $\alpha\notin\overline{\varphi}(v)$ (a common situation since $U$ is elementary in their Lemma 3.2). In this case, $v$ is no longer an endpoint of the $(\alpha,\alpha_{2})$-path containing $R$ under $\varphi^{*}=\varphi/(G-U,\alpha_{1},\alpha)$. Consequently, their assumption that $w\in V(T_{n}(x))$ fails (see paragraph 6 after Claim 1 in their Lemma 3.2, which corresponds to (3) in the proof of our Lemma 4.3).
Furthermore, when applying their Lemma 3.2 in the proof of their Lemma 3.3 (which is analogous to our Lemma 4.1 together with Claim 5.3), the argument fails when applied to $P_{x}$ in paragraph 6 of their Lemma 3.3. This corresponds precisely to Case II in the proof of our Lemma 4.1, where Lemma 4.6 is essential to resolve the issue. Consequently, our proof of Lemma 4.1 must be more technical in order to properly apply Lemma 4.6.
In fact, Lemma 4.6 is essential to our approach. Its proof requires extending both the techniques and the algorithm (see the paragraph after $(f_{i})$ in the proof of Lemma 4.5) to their full strength. In contrast, they use this algorithm only once in their proof of Lemma 3.2 (see the algorithm introduced before their Claim 2), which explains why their paper is shorter but also why their argument is incomplete.
@@
(A2.5) Let $n$ be a nonnegative integer. Suppose (A1) and (A2) hold for all ETTs with at most $n$ rungs, and $(A2.5)$ holds for all ETTs with at most (n1)(n-1) rungs. Let $T_{n+1}$ be a closed ETT with ladder $T_{0}\subset\dots\subset T_{n}\subset T_{n+1}$. Suppose $T_{n+1}$ could be extended further to $T_{n+2}$ with $F_{n+1}=\{f_{n+1}\}$, $S_{n+1}=\{\gamma_{n+1},\delta_{n+1}\}$, and initial index $h$. If $\varphi^{\prime}$ is a $(T_{n+1},D_{n+1},\varphi)$-stable coloring such that $T_{h}$ has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\varphi^{\prime}$ as under $\varphi$, then $\varphi^{\prime}$ is a $T_{n+1}+f_{n+1}$ mod $\varphi$ coloring.
@@
Let $T_{n+1}$, $F_{n+1}$, $S_{n+1}$, $h$, and $\varphi^{\prime}$ be as described above. We first consider the case when extension type $\Theta_{n+1}=\operatorname{IE}$. Then $T_{n+1}$ is $\operatorname{RE}$ finished, $h=n+1$, $\delta_{n+1}$ is a defective color of $T_{n+1}$, and $f_{n}$ belongs to a $(\gamma_{n+1},\delta_{n+1})$-ear of $T_{n+1}$ under $\varphi$. Moreover, since $T_{n+1}$ is elementary by (A1), $T_{n+1}$ has exactly one $(\gamma_{n+1},\delta_{n+1})$-exit vertex under $\varphi$ by (A2). Since $\varphi^{\prime}$ is $(T_{n+1},D_{n+1},\varphi)$-stable, $T_{n+1}$ is closed under $\varphi^{\prime}$ and $\varphi^{\prime}$ is $(T_{n+1},D_{n},\varphi)$-stable. So $\varphi^{\prime}$ is a $T_{n+1}$ mod $\varphi$ coloring by (A2), and therefore
(1) $T_{n+1}$ is a closed ETT under $\varphi^{\prime}$ with the same ladder, $F_{i}$, $S_{i}$, and $\Theta_{i}$ for $0\leq i\leq n$, and $T_{n+1}$ is still $\operatorname{RE}$ finished under $\varphi^{\prime}$.
Recall that $f_{n+1}$ belongs to a $(\gamma_{n+1},\delta_{n+1})$-ear of $T_{n+1}$ under $\varphi$. Since $\varphi^{\prime}$ is a $(T_{n+1},D_{n+1},\varphi)$-stable coloring such that $T_{n+1}$ has the same $(\gamma_{n+1},\delta_{n+1})$-exit vertex under $\varphi^{\prime}$ as $\varphi$, the connecting edge $f_{n+1}$ still belongs to a $(\gamma_{n+1},\delta_{n+1})$-ear of $T_{n+1}$ under $\varphi^{\prime}$. Since $T_{n+1}$ is still $\operatorname{RE}$ finished under $\varphi^{\prime}$ by (1), we may still choose $F_{n+1}=\{f_{n+1}\}$ and $S_{n+1}=\{\gamma_{n+1},\delta_{n+1}\}$ to extend $T_{n+1}$ by an $\operatorname{IE}$ extension under $\varphi^{\prime}$, so the coloring $\varphi^{\prime}$ is a $T_{n+1}+f_{n+1}$ mod $\varphi$ coloring following (1).
We then consider the case when extension type $\Theta_{n+1}=\operatorname{RE}$ with initial index $h\leq n$. In this case, $n\geq 1$ and $h\geq 1$. Since $T_{h}$ is elementary by (A1), $T_{h}$ has exactly one $(\gamma_{h},\delta_{h})$-exit vertex under $\varphi$ by (A2). By the assumption on $\varphi^{\prime}$, the ETT $T_{h}$ has the same unique $(\gamma_{h},\delta_{h})$-exit vertex under $\varphi^{\prime}$ as under $\varphi$. Note that in this case $\varphi^{\prime}$ is also a $(T_{n},D_{n},\varphi)$-stable coloring, we see that $\varphi^{\prime}$ is a $T_{n}+f_{n}$ mod $\varphi$ coloring by (A2.5) with $T_{n}$ that has (n1)(n-1)-rungs in place of $T_{n+1}$. As $\varphi^{\prime}$ is also $(T_{n+1},D_{n+1},\varphi)$-stable,
(2) $T_{n+1}$ is a closed ETT under $\varphi^{\prime}$ with the same ladder, $F_{i}$, $S_{i}$, and $\Theta_{i}$ for $0\leq i\leq n$.
Furthermore, $f_{n+1}$ is still colored by $\gamma_{h}=\gamma_{n+1}$. Because $\Theta_{n+1}=\operatorname{RE}$, the connecting edge $f_{n+1}$ belongs to a $(\gamma_{h},\delta_{h})$-ear $O$ of $T_{h}$ such that $O$ contains a path $L$ connecting $f_{n+1}$ and $V(T_{h})$ with $V(L)\subseteq V(T_{n+1})$ under $\varphi$. Since $\varphi^{\prime}$ is $(T_{n+1},D_{n+1},\varphi)$-stable, $L$ remains a path contained entirely in either a $(\gamma_{h},\delta_{h})$-ear or a $(\gamma_{h},\delta_{h})$-exit path of $T_{h}$ under $\varphi^{\prime}$. Since $T_{h}$ has the same unique $(\gamma_{h},\delta_{h})$-exit vertex under $\varphi^{\prime}$ as under $\varphi$, the path $L$ must be contained entirely in a $(\gamma_{h},\delta_{h})$-ear $O^{*}$ of $T_{h}$ under $\varphi^{\prime}$ that connects $f_{n+1}$ to $V(T_{h})$ (here $O^{*}$ might be different from $O$). So we may still choose $F_{n+1}=\{f_{n+1}\}$ and $S_{n+1}=\{\gamma_{n+1},\delta_{n+1}\}$ to extend $T_{n+1}$ by an $\operatorname{RE}$ extension under $\varphi^{\prime}$, and therefore the coloring $\varphi^{\prime}$ is a $T_{n+1}+f_{n+1}$ mod $\varphi$ coloring following (2). ∎
@@
(A3) Let $n$ be a nonnegative integer. Suppose (A1), (A2), and (A2.5) hold for all ETTs with at most $n$ rungs, and (A3) holds for all ETTs with at most (n1)(n-1) rungs. Let $T_{n+1}$ be a closed ETT with ladder $T_{0}\subset\dots\subset T_{n}\subset T_{n+1}$. Assume $T_{n+1}$ is $\operatorname{RE}$ finished, $\delta$ is a defective color of $T_{n+1}$ under $\varphi$, and $\gamma\in\overline{\varphi}(T_{n+1})$. Let $\alpha,\beta$ be two colors such that $\alpha\in\overline{\varphi}(T_{n+1})\setminus\{\gamma\}$ and $\beta$ is arbitrary. For any $(\alpha,\beta)$-path $P$, if $P$ does not intersect $T_{n+1}$, then $T_{n+1}$ has the same $(\gamma,\delta)$-exit vertex under $\varphi/P$ as under $\varphi$.